OpenStudy (anonymous):

Fan and Medal!! The ages of three brothers are consecutive integers with the sum of 96. a)Write an expression to determinate the ages of the brothers. b)How old are the three brothers?

OpenStudy (zaraxx):

i'll help just give me some time to type

OpenStudy (zaraxx):

bro #1= x bro #2= x+1 bro#3= x+2 u want all 3 to add up to 96 so x+x+1+x+2=96 combine like terms 3x+3=96 subtract 3 from both sides 3x=96 divide both sides by 3 31 is the answer therefore the next brothers ages will be 32 and 33 to check, add 31+32+33 does this make sense?

OpenStudy (dm91):

31+32+33= 97 so your answer would be 31, 32, and 33

OpenStudy (zaraxx):

yes it would be 96 so those are u answers!

OpenStudy (anonymous):

thanks guys ur the best u my homie

OpenStudy (dm91):

welcome